OpenStudy (anonymous):

HELP! Im studying for finals!!! find y=f^-1(x) if f(x)=5-2x^2, x>0

OpenStudy (anonymous):

\[f^-1(x)=\sqrt{x-5}/-2\] is this the right answer?

jimthompson5910 (jim_thompson5910):

the -2 should be under the root, you should have \[\Large f^{-1}(x) = \sqrt{\frac{x-5}{-2}}\] which simplifies to \[\Large f^{-1}(x) = \sqrt{\frac{5-x}{2}}\]

jimthompson5910 (jim_thompson5910):

the reason why is because you apply the square root after you divide by -2